OpenStudy (anonymous):

How much energy does a 50-watt light bulb use compared to a 100-watt light bulb if both are shining for the same length of time? Explain your answer.

OpenStudy (ujjwal):

Energy consumed = power *time Energy used by 50 watt bulb = 50 *t Energy used by 100 watt bulb = 100 *t where t is the time for which they are shining from above it is clear that the energy used by 100 watt bulb is twice the energy used by 50 watt bulb.

OpenStudy (anonymous):

Thank you very much....I was beginning to get a headache trying to figure the answer out.